On May 31 of the current year, the assets and liabilities of Riser, Inc. are as follows: Cash $14,800; Accounts Receivable, $6,950; Supplies, $500; Equipment, $11,550; Accounts Payable, $8,850. What is the amount of equity as of May 31 of the current year?
Answer #1

Cash + Accounts Receivable + Supplies + Equipment = Accounts Payable + Equity

14,800+6,950+500+11,550 = 8,850+Equity

33,800 = 8,850+Equity

Therefore, Equity = 33,800-8,850 = $24,950